# Pulse-PPG: An Open-Source Field-Trained PPG Foundation Model for Wearable Applications Across Lab and Field Settings

### v1

@@ Abstract.
Photoplethysmography (PPG)-based foundation models are gaining traction due to the widespread use of PPG in biosignal monitoring and their potential to generalize across diverse health applications. In this paper, we introduce Pulse-PPG, the first open-source1 PPG foundation model trained exclusively on raw PPG data collected over a 100-day field study with 120 participants. Existing PPG foundation models are either open-source but trained on clinical data or closed-source, limiting their applicability in real-world settings. We evaluate Pulse-PPG across multiple datasets and downstream tasks, comparing its performance against a state-of-the-art foundation model trained on clinical data. Our results demonstrate that Pulse-PPG, trained on uncurated field data, exhibits superior generalization across clinical and mobile health applications in both lab and field settings. This suggests that exposure to real-world variability enables the model to learn fine-grained representations, making it more adaptable across tasks. Furthermore, pre-training on field data surprisingly outperforms its pre-training on clinical data in many tasks, reinforcing the importance of training on real-world, diverse datasets. To encourage further advancements in robust foundation models leveraging field data, we plan to release Pulse-PPG, providing researchers with a powerful resource for developing more generalizable PPG-based models.
@@
Both authors contributed equally to this research.
copyright: none
@@ 1. Introduction
Figure 1. Overview of our Pulse-PPG Foundation Model. Our model is trained on wearable field PPG with relative contrastive learning, based on the relative distances captured in our learned motif-based distance function. This model demonstrates strong performance on a wide variety of tasks across wearable field, wearable lab, and clinical settings.
The emergence of Foundation Models (FM) offers a new opportunity to address these challenges and accelerate our progress. In other domains such as natural language processing and computer vision, the foundation model paradigm has transformed the development of machine learning solutions to real-world problems. The key property of an FM is that it is pre-trained on a large-scale dataset to ensure that the resulting feature representation encompasses all of the complexity of the data domain, which is then validated by demonstrating that the FM can solve multiple downstream tasks without additional representation learning. In computer vision, the field has moved away from collecting individual special-purpose datasets and training task-specific models to leveraging existing foundation model representations, such as DINOv2 (Oquab et al., 2023) in solving a variety of perceptual tasks. A key enabling property is that, while the training datasets are private and not publicly available, the model weights are released to the research community, enabling everyone to benefit from its powerful representation. This transition in utilizing publicly available FMs has not yet occurred for mHealth, and it is a crucial next step.
@@
In this paper, we present the first open-source [^1] Model weights and code will be released under the MIT license (Saltzer, 2020) upon acceptance of our work. foundation model for wearable PPG signals, Pulse-PPG, which can be seen in Figure 1. We pre-train our model on the raw PPG signals from a large-scale wearable field dataset, composed of 21 billion data points collected from 120 participants, organized in 822,247 unique 4-minute PPG segments. We avoid filtering our signals because such real-world noise patterns offer meaningful contextual cues on a given individual’s motion or environment, which we show, enables better performance on wearable field, lab and clinical downstream tasks, compared to training on cleaner clinical data. In order to use our foundation model for the field setting, we utilize a learnable motif-based distance function (Xu et al., 2024a), as well a relative contrastive learning loss (Xu et al., 2024b) in order to extract subtle but semantically meaningful patterns from the data. To this end, our model achieves state-of-the-art performance, substantially outperforming another state-of-the-art open sourced PPG foundation model on 10/11 of our downstream evaluation tasks that span 5 different datasets, across the wearable field, wearable lab, and clinical PPG domains.
With the weights publicly available, researchers will be able to build upon our work for their own specialized tasks and can bypass the costly and time-consuming process of training from scratch, significantly lowering the barrier to entry for those without access to extensive computational resources. Moreover, our model can serve as an embedding function, enabling researchers to extract meaningful physiological representations from raw PPG data. Beyond these, the availability of our model weights will encourage reproducibility and benchmarking across datasets and settings. Researchers can compare their methods against a common baseline, accelerating research progress.
@@
    Pulse-PPG is the first open-source1 PPG foundation model trained entirely on raw, uncurated PPG data collected from a longitudinal field study involving 120 subjects from diverse backgrounds over 100 days.
@@
    We achieve state-of-the-art performance across a wide set of downstream datasets and tasks for wearable applications within field and lab settings, as well as clinical applications.
@@
    We show that for two different foundation model approaches, training on field data achieves better performance than a lab-trained model on wearable lab downstream tasks, highlighting the field-to-lab generalizability of the models, with additional potential implications for clinical settings.
@@ 2.1. Foundation Models for PPG signals
There have been some recent work on PPG-specific foundation models, but none of them are open and designed for wearable PPG signals (i.e., collected from a smartwatch). The closest work with open model availability is PaPaGei, an open-source PPG foundation model that improves classification and regression performance across 14 tasks (Pillai et al., 2024), including some wearable lab tasks. However, it is exclusively trained on clean clinical PPG signals, and was not evaluated on field data in (Pillai et al., 2024). We show (Section 5.2) that training on this clinical data negatively impacted its generalization performance on wearable tasks involving field-collected data. The most relevant work for wearable PPGs is a foundation model pre-trained on a large-scale wearable field PPG dataset and presented in (Abbaspourazad et al., 2023). However, the model is closed-source and its training datasets are private, restricting accessibility for the research community. Similalry, SiamQuality is a foundation model pre-trained on clinical data that demonstrates generalizability on wearable lab and clinical datasets, but the model remains private (Ding et al., 2024).
Additional related works addressed tasks and goals which are adjacent to this paper. REGLE employs autoencoders to extract disentangled embeddings from clinical PPGs, focusing on their applicability to genetic and disease prediction tasks (Yun et al., 2024). TS2TC introduces a generative self-supervised learning framework for clinical PPG data but does not provide pre-trained weights and has been evaluated exclusively on clinical PPG regression tasks (Zhang et al., 2024b). PPG-PT demonstrates the utility of pre-training, but does not meet all of the requirements of a foundation model, as the model was only evaluated on one clinical PPG-specific downstream task (Davies et al., 2024).
While these prior works have made significant strides in PPG representation learning, they either lack open-source accessibility, generalizability across tasks, or applicability to field-collected wearable data, highlighting a critical gap that our work aims to address. With the release of our foundation model’s weights, ours will be the first open-source1 PPG model trained exclusively on field-collected data.
@@ 2.2. Foundation Models in General
Our work is inspired by prior successes in developing foundation models for other data modalities. These works have shown that models pre-trained on large-scale datasets can capture underlying data representations and are highly adaptable for various downstream applications (Bommasani et al., 2021). These models have had substantial impact in fields like Natural Language Processing and Computer Vision. In the language domain, models such as BERT (Devlin et al., 2018) and the Generative Pretrained Transformer (GPT) (Radford et al., 2018) have demonstrated remarkable generalizability across a diverse range of natural language processing tasks, catalyzing the development of large language models (i.e. ChatGPT (Zhou et al., 2024)). The vision domain has seen similar breakthroughs with models like CLIP (Radford et al., 2021) and the Segment Anything Model (Kirillov et al., 2023) that are trained on large-scale data and exhibit strong performance on a broad set of downstream tasks. These works have demonstrated that pre-training on large-scale data can yield effective feature representations and serve as the motivation for this work.
3. Designing Pulse-PPG Methodology
@@ 3.1. Motivation for Pre-training on Field PPG data
Traditionally, prior work has focused on training models using clean, lab-collected PPG data and evaluating them on mHealth datasets (Pillai et al., 2024). This approach is largely due to the significant challenges associated with utilizing real-world field data, which is inherently noisy and variable. However, we hypothesize that by designing a methodology capable of effectively learning patterns from field data, we can develop representations that are more robust, generalize better, and achieve superior performance on downstream mHealth tasks, such as instantaneous heart rate prediction or activity recognition. In this way, the noise present from the wearable field sensor becomes a benefit, as it could introduce additional contextual information of a person’s current state and environment. Importantly, pre-training on mHealth data and evaluating on mHealth data reduces domain shift, as the model is exposed to the same types of noise, variability, and environmental factors during both training and evaluation.
@@
3.2. Designing the Pre-training Task for the Pulse-PPG Foundation Model
@@
To build a robust foundation model, we must design a pre-training task that not only captures high-level, obvious semantic features—such as heart rate changes—but also uncovers subtle, latent patterns in the data, such as variations in pulse waveform morphology or transient physiological responses. PPG signals, which measure blood volume changes, encode a wealth of physiological information. However, this information is often obscured by noise, particularly in mHealth settings. Traditional approaches address noise by normalizing or aggregating data into statistical features and training models on these features. While this reduces the impact of noise, it often discards fine-grained details that are critical for accurate inference or discovering hidden clusters and trends in the data.
@@
To overcome these challenges, we introduce Pulse-PPG, a method specifically designed to address three key issues: 1) the development of a domain-specific pre-training task that leverages PPG-specific domain knowledge; 2) exploiting the inherent noise of real-world, mHealth PPG data; and 3) the ability to capture subtle, yet meaningful, patterns in the data that are essential for ensuring robustness and generalizability across diverse and dynamic real-world settings.
3.2.1. PPG-Specific Pre-training Task:
@@
A classic self-supervised objective for time-series foundation models is contrastive learning, where augmentations (e.g., cropping, scaling, jittering, shuffling) create distinct but semantically similar positive pairs (Meng et al., 2023; Yue et al., 2022; Zhang et al., 2022; Eldele et al., 2023; Woo et al., 2022; Yang and Hong, 2022; Yang et al., 2022; Ozyurt et al., 2022; Lee et al., 2022). However, PPG signals face unique challenges: (1) time-series lack image-like invariances (e.g., rotation), and (2) augmentations can distort PPG semantics. For example, scaling alters amplitude (critical for blood volume inference) and shuffling disrupts temporal dependencies (key for heart rate variability). Alternatively, subject-aware contrastive learning avoids augmentations by leveraging long-term recordings from individual subjects (Abbaspourazad et al., 2024). While especially effective for biosignals (Jeong et al., 2024), existing methods rely on naive temporal sampling (Tonekaboni et al., 2021) or supplemental context (e.g. event labels) (Jeong et al., 2023). To address this, we propose directly exploiting raw PPG semantics in our large-scale unlabeled dataset, eliminating reliance on augmentations or auxiliary data. To unlock the potential of our large-scale unlabeled wearable field PPG dataset, we need to design a methodology that utilizes semantic information from only the raw PPG signals.
@@
A critical characteristic of raw PPG signals lies in their origins in the cardiopulmonary system. PPG signals are pulsative: quasiperiodic and composed of repeating motifs that correspond to individual cardiac cycles. We define motifs abstractly as the short temporal shapes within the overall PPG morphology that reflect specific semantic information, such as the first upward slope in a specific PPG beat that represents systolic rise time. These motifs encode critical physiological information, such as heart rate variability, blood pressure changes, and vascular dynamics.
@@
Therefore, we design our pre-training task to learn differences between PPG instances, based upon the differences in their motifs. In this way, we utilize a distance function, $d(\textbf{X}_{\text{anchor}},\textbf{X}_{\text{cand}})$ that compares two PPG sequences, $\textbf{X}_{\text{anchor}}\in\mathbb{R}^{T}$ and $\textbf{X}_{\text{cand}}\in\mathbb{R}^{T}$ by finding the motif in $\textbf{X}_{\text{cand}}$ that best matches a given motif $\textbf{X}_{\text{anchor}}$. This idea can be seen below in Equation 1, with the ”matching” function given by a distance function $d_{m}$.
@@
3.2.2. Exploiting the Noise of Real-World, mHealth PPG data:
Many prior works that utilize morphology information from raw PPG signals for their task-specific application, such as for blood pressure prediction (Salah et al., 2022; Shin and Min, 2017) or stress detection (Ahmed et al., 2022; Rinkevičius et al., 2019), will first utilize a peak detection algorithm to segment out individual PPG beats that correspond to as single cardiac cycle (Haddad et al., 2020; Kavsaoğlu et al., 2016; Fischer et al., 2016). After segmentation, specific temporal features within the beat can be investigated, such as the dicrotic notch or the systolic rise time. However, these methods are evaluated exclusively on clinical or lab-collected PPG data, where signal quality is controlled and high.
Exploiting PPG morphological information is significantly more challenging in real-world wearable field PPG data, where persistent noise and motion artifacts, make beat-to-beat segmentation and morphological feature extraction infeasible. As shown in Figure 2, the sharp decline in PPG signal quality—ubiquitous in mobile health settings—severely undermines the reliability of traditional segmentation and feature isolation methods.

### v2

@@ Abstract.
Photoplethysmography (PPG)-based foundation models are gaining traction due to the widespread use of PPG in biosignal monitoring and their potential to track diverse health indicators. In this paper, we introduce Pulse-PPG, an open-source PPG foundation model trained exclusively on raw PPG data collected over a 100-day field study with 120 participants. Existing open-source PPG foundation models are trained on clinical data, and those trained on field data are closed source, limiting their applicability in real-world settings. Extensive evaluations demonstrate that Pulse-PPG, trained on uncurated field data, exhibits superior generalization and performance across clinical and mobile health applications in both lab and field settings, when compared with state-of-the-art PPG foundation models trained on clinical data. Exposure to real-world variability in field-collected PPG data enables Pulse-PPG to learn more robust representations. Furthermore, pre-training Pulse-PPG on field data outperforms its own pre-training on clinical data in many tasks, reinforcing the importance of training on real-world datasets. To encourage further advancements in robust PPG modeling, we have open-sourced2our Pulse-PPG model, providing researchers with a valuable resource for developing the next generation of task-specific PPG-based models.
@@
$\bm{\dagger}$Both co-first authors contributed equally to this research.
copyright: acmlicensedjournal: IMWUTjournalyear: 2025journalvolume: 9journalnumber: 3article: 126publicationmonth: 9doi: 10.1145/3749494ccs: Human-centered computing Ubiquitous and mobile computingccs: Computing methodologies Machine learning
@@ 1. Introduction
Figure 1. Overview of our open-sourced222 Pulse-PPG model weights and code repository are available at https://github.com/maxxu05/pulseppg. Pulse-PPG Foundation Model. Our model is trained on wearable field PPG with relative contrastive learning, based on the relative distances captured in our learned motif-based distance function. This model demonstrates strong performance on a wide variety of tasks across wearable field, wearable lab, and clinical settings.
The emergence of Foundation Models (FM) offers a new opportunity to address these challenges and accelerate our progress. In other domains such as natural language processing and computer vision, the foundation model paradigm has transformed the development of machine learning solutions to real-world problems (App, 2025; Met, 2025). The key property of an FM is that it is pre-trained on a large-scale dataset to ensure that the resulting feature representation encompasses all of the complexity of the data domain, which is then validated by demonstrating that the FM can solve multiple downstream tasks without additional representation learning. In computer vision, the field has moved away from collecting individual special-purpose datasets and training task-specific models to leveraging existing foundation model representations, such as DINOv2 (Oquab et al., 2023) in solving a variety of perceptual tasks. A key enabling property is that while the training datasets are private and not publicly available, the model weights are released to the research community, enabling everyone to benefit from its powerful representation. This transition in utilizing publicly available FMs has not yet occurred for mHealth and is a crucial next step.
@@
In this paper, we present an open-source2 foundation model for wearable field PPG signals, Pulse-PPG, which can be seen in Figure 1. We pre-train our model on the raw PPG signals from a large-scale wearable field dataset, composed of approximately 200 million seconds of raw PPG data collected from 120 participants. We avoid filtering our signals because such real-world noise patterns offer meaningful contextual cues on a given individual’s motion or environment, which we show enables better performance on wearable field, lab, and clinical downstream tasks, compared to training on cleaner clinical data. To address the unique challenges of field-collected PPG data, we adapted a motif-based relative contrastive learning framework (Xu et al., 2024a, c) to flexibly handle the temporal and noise characteristics of the field PPG signals and extract subtle but semantically meaningful patterns. Our empirical evaluation demonstrates that this approach achieves state-of-the-art performance, outperforming a recent PPG foundation model on 10/11 of our downstream evaluation tasks across five different datasets, covering wearable data from the field, wearable data from the lab, and clinical PPG domains.
With the weights publicly available, researchers can build upon our work for their own specialized tasks and bypass the costly and time-consuming process of training from scratch, significantly lowering the barrier to entry for those without access to extensive computational resources. Moreover, our model can serve as an embedding function, enabling researchers to extract meaningful physiological representations from raw PPG data. Beyond these, the availability of our model weights will encourage reproducibility and benchmarking across datasets and settings. Researchers can compare their methods against a common baseline, accelerating research progress.
@@
    We present Pulse-PPG, the best-performing open-source2 PPG foundation model trained entirely on raw, uncurated PPG data collected from a 100-day field study involving 120 subjects from diverse backgrounds.
@@
    We show that training a motif-based representation learning framework with relative contrastive learning outperforms the current state-of-the-art open-source PPG foundation model (Pillai et al., 2024) across a wide set of downstream datasets and tasks for wearable field and clinical applications. We also show that our Pulse-PPG model with 28.5 million parameters outperforms substantially more powerful general-purpose time-series models—Chronos with 200 million parameters (Ansari et al., 2024) and MOMENT with 385 million parameters (Goswami et al., 2024).
@@
    We show that training on field data achieves better performance for two different foundation model approaches than a lab-trained model on wearable lab downstream tasks, highlighting the field-to-lab generalizability of the models, with additional potential implications for clinical settings.
@@ 2.1. Foundation Models for PPG signals
There have been some recent work on PPG-specific foundation models, but none of them are open and designed for wearable PPG signals (i.e., collected from a smartwatch). The closest work with open model availability is PaPaGei, an open-source PPG foundation model that improves classification and regression performance across 14 tasks (Pillai et al., 2024), including some wearable lab tasks. However, it is exclusively trained on clean clinical PPG signals, and was not evaluated on field data (Pillai et al., 2024). We show (Section 6.1) that training on this clinical data negatively impacted its generalization performance on wearable tasks involving field-collected data. The most relevant work for wearable PPGs is a foundation model pre-trained on a large-scale wearable field PPG dataset and presented in (Abbaspourazad et al., 2023). However, the model is closed-source and its training datasets are private, restricting accessibility for the research community. Similarly, SiamQuality is a foundation model pre-trained on clinical data that demonstrates generalizability on wearable lab and clinical datasets, but the model remains private (Ding et al., 2024).
Additional related works addressed tasks and goals that are adjacent to this paper. REGLE employs autoencoders to extract disentangled embeddings from clinical PPGs, focusing on their applicability to genetic and disease prediction tasks (Yun et al., 2024). TS2TC introduces a generative self-supervised learning framework for clinical PPG data but does not provide pre-trained weights and has been evaluated exclusively on clinical PPG regression tasks (Zhang et al., 2024b). PPG-PT demonstrates the utility of pre-training, but does not meet all of the requirements of a foundation model, as the model was only evaluated on one clinical PPG-specific downstream task (Davies et al., 2024).
While these prior works have made significant strides in PPG representation learning, they either lack open-source accessibility, generalizability across tasks, or applicability to field-collected wearable data, highlighting a critical gap that our work aims to address. With the release of our foundation model’s weights, ours will be the first open-source2 PPG model trained exclusively on field-collected data.
@@ 2.2. Foundation Models in General
Our work is inspired by prior successes in developing foundation models for other data modalities. These works have shown that models pre-trained in large-scale datasets can capture underlying data representations and are highly adaptable for various downstream applications (Bommasani et al., 2021). These models have had a substantial impact in fields like Natural Language Processing and Computer Vision. In the language domain, models such as BERT (Devlin et al., 2018) and the Generative Pretrained Transformer (GPT) (Radford et al., 2018) have demonstrated remarkable generalizability across a diverse range of natural language processing tasks, catalyzing the development of large language models (i.e., ChatGPT (Zhou et al., 2024)). The vision domain has seen similar breakthroughs with models like CLIP (Radford et al., 2021) and the Segment Anything Model (Kirillov et al., 2023) that are trained on large-scale data and exhibit strong performance on a broad set of downstream tasks. These works have demonstrated that pre-training on large-scale data can yield effective feature representations and serve as the motivation for this work.
3. Pulse-PPG Methodology
@@ 3.1. Motivation for Pre-training on Field PPG data
Traditionally, prior work has focused on training models using clean, lab-collected PPG data and evaluating them on mHealth datasets (Pillai et al., 2024). This approach is largely due to the significant challenges associated with utilizing real-world field data, which is inherently noisy and variable. However, we hypothesize that by designing a methodology capable of effectively learning patterns from field data, we can develop more robust representations, generalize better, and achieve superior performance on downstream mHealth tasks, such as instantaneous heart rate prediction or activity recognition. In this way, the noise present from the wearable field sensor becomes a benefit, as it could introduce additional contextual information of a person’s current state and environment. Importantly, pre-training on mHealth data and evaluating on mHealth data reduces domain shift, as the model is exposed to the same types of noise, variability, and environmental factors during both training and evaluation.
@@
3.2. Learning Robust PPG Representations from Field Data
@@
To build a robust foundation model, we must design a pre-training task that not only captures high-level, obvious semantic features, such as heart rate changes, but also uncovers subtle, latent patterns in the data, such as variations in pulse waveform morphology or transient physiological responses. PPG signals, which measure blood volume changes, encode a wealth of physiological information. However, this information is often obscured by noise, particularly in the field. Traditional approaches address noise by normalizing or aggregating data into statistical features and training models on these features. While this reduces the impact of noise, it often discards fine-grained details that are critical for accurate inference or discovering hidden clusters and trends in the data.
@@
To overcome these challenges, we introduce Pulse-PPG, a method specifically designed to address three key issues: 1) the development of a domain-specific pre-training task that leverages PPG-specific domain knowledge; 2) exploiting the inherent noise of real-world PPG data; and 3) the ability to capture subtle, yet meaningful, patterns in the data that are essential for ensuring generalizability across diverse and dynamic real-world settings.
3.2.1. PPG-Specific Pre-training Task
@@
A classic self-supervised objective for time-series foundation models is contrastive learning, where augmentations (e.g., cropping, scaling, jittering, shuffling) create distinct but semantically similar positive pairs (Meng et al., 2023; Yue et al., 2022; Zhang et al., 2022; Eldele et al., 2023; Woo et al., 2022; Yang and Hong, 2022; Yang et al., 2022; Ozyurt et al., 2022; Lee et al., 2022). However, PPG signals face unique challenges: (1) time-series lack image-like invariances (e.g., rotation), and (2) augmentations can distort PPG semantics. For example, scaling alters amplitude (critical for blood volume inference) and shuffling disrupts temporal dependencies (key for heart rate variability). Alternatively, subject-aware contrastive learning avoids augmentations by leveraging long-term recordings from individual subjects (Abbaspourazad et al., 2023). While especially effective for biosignals (Jeong et al., 2024), existing methods rely on naive temporal sampling (Tonekaboni et al., 2021) or supplemental context (e.g. event labels) (Jeong et al., 2023). To address this, we propose directly exploiting raw PPG semantics in our large-scale unlabeled dataset, eliminating reliance on augmentations or auxiliary data. To unlock the potential of our large-scale, unlabeled wearable field PPG dataset, we need a methodology that utilizes semantic information from only the raw PPG signals.
@@
A critical characteristic of raw PPG signals lies in their origins in the cardiopulmonary system. PPG signals are pulsative: quasiperiodic and composed of repeating motifs that correspond to individual cardiac cycles (Allen, 2007; Elgendi, 2012). We define motifs abstractly as the short temporal shapes within the overall PPG morphology that reflect specific semantic information, such as the first upward slope in a specific PPG beat that represents systolic rise time. These motifs encode critical physiological information, such as heart rate variability, blood pressure changes, and vascular dynamics.
@@
Therefore, we propose to utilize our pre-training task to learn differences between PPG instances, based upon the differences in their motifs, drawing from Xu et al. (2024b). In this way, the distance function, $d(\textbf{X}_{\text{anchor}},\textbf{X}_{\text{cand}})$ can compare two PPG sequences, $\textbf{X}_{\text{anchor}}\in\mathbb{R}^{T}$ and $\textbf{X}_{\text{cand}}\in\mathbb{R}^{T}$ by finding the motif in $\textbf{X}_{\text{cand}}$ that best matches a given motif $\textbf{X}_{\text{anchor}}$. This idea can be seen below in Equation 1, with the ”matching” function given by a distance function $d_{m}$.
@@
3.2.2. Exploiting the Noise of Real-World, mHealth PPG data
Many prior works that utilize morphology information from raw PPG signals for their task-specific application, such as for blood pressure prediction (Salah et al., 2022; Shin and Min, 2017) or stress detection (Ahmed et al., 2022; Rinkevičius et al., 2019), will first utilize a peak detection algorithm to segment individual PPG beats that correspond to as single cardiac cycle (Haddad et al., 2020; Kavsaoğlu et al., 2016; Fischer et al., 2016). After segmentation, specific temporal features within the beat can be investigated, such as the dicrotic notch or the systolic rise time. However, these methods are evaluated exclusively on clinical or lab-collected PPG data, where signal quality is controlled and high.
Exploiting PPG morphological information is significantly more challenging in real-world wearable field PPG data, where persistent noise and motion artifacts make beat-to-beat segmentation and morphological feature extraction infeasible. As shown in Figure 2, the sharp decline in PPG signal quality, ubiquitous in mobile health settings, severely undermines the reliability of traditional segmentation and feature isolation methods.